# Beyond Fact Retrieval: Episodic Memory for RAG with Generative Semantic Workspaces

### v1

@@ Abstract
Large Language Models (LLMs) face fundamental challenges in long-context reasoning: many documents exceed their finite context windows, while performance on texts that do fit degrades with sequence length, necessitating their augmentation with external memory frameworks. Current solutions, which have evolved from retrieval using semantic embeddings to more sophisticated structured knowledge graphs representations for improved sense-making and associativity, are tailored for fact-based retrieval and fail to build the space-time-anchored narrative representations required for tracking entities through episodic events. To bridge this gap, we propose the Generative Semantic Workspace (GSW), a neuro-inspired generative memory framework that builds structured, interpretable representations of evolving situations, enabling LLMs to reason over evolving roles, actions, and spatiotemporal contexts. Our framework comprises an Operator, which maps incoming observations to intermediate semantic structures, and a Reconciler, which integrates these into a persistent workspace that enforces temporal, spatial, and logical coherence. On the Episodic Memory Benchmark (EpBench) (huet_episodic_2025) comprising corpora ranging from 100k to 1M tokens in length, GSW outperforms existing RAG based baselines by up to 20%. Furthermore, GSW is highly efficient, reducing query-time context tokens by 51% compared to the next most token-efficient baseline, reducing inference time costs considerably. More broadly, GSW offers a concrete blueprint for endowing LLMs with human-like episodic memory, paving the way for more capable agents that can reason over long horizons.
Introduction
@@
Large Language Models (LLMs) have transformed natural language understanding, but their ability to reason over long contexts is still limited by finite input windows. Even with token limits in the millions, large document collections can easily exceed these bounds. Performance can also degrade with context length due to phenomena like “context rot” and “lost-in-the-middle” effects (liu_lost_2023; hong2025context). A common workaround is Retrieval-Augmented Generation (RAG), which supplements the LLM’s input with only the most relevant retrieved content at query time. Standard RAG pipelines split documents into smaller chunks, encode them into dense embeddings, and retrieve the top-matching chunks based on semantic similarity to the query—allowing the LLM to focus on a relevant subset of the corpus during inference.
@@
A key limitation of standard RAG methods is that each text chunk is embedded independently, which can lead to incomplete retrieval when a query depends on information spread across multiple chunks. Because similarity scores are computed in isolation, essential context may be missed. To address this, more recent approaches have adopted structured representations — such as knowledge graphs — that explicitly model relationships between chunks across the corpus. At query time, these graphs are traversed or queried to retrieve semantically connected chunks, enabling LLMs to perform more effective multi-hop reasoning and question answering (gutierrez_hipporag_2025; JimenezGutierrez2025HippoRAG; Edge2025GraphRAG; Guo2024LightRAG).
@@
These methods have primarily been evaluated on fact-rich documents such as Wikipedia pages (yang2018hotpotqa; xanh2020_2wikimultihop; trivedi2022musiquemultihopquestionssinglehop). Yet the vast majority of texts that LLMs encounter are not lists of facts but narratives of evolving real-world situations. Crime reports, political briefings, corporate filings, legislative records, war dispatches, and multi-day news coverage all describe actors (people, organizations, nations) that adopt roles (suspect, regulator, bidder, combatant) and transition through states (arrested → arraigned → released; startup → unicorn → acquired) while interacting across space and time.
@@
We contend that reasoning over such documents would be much more accurate and energy efficient, if one indexed the documents in terms of an internal world model— a structured representation that keeps track of who is involved, what was done, where and when events occur, how roles change, and what consequences follow. Indeed, to achieve such a goal, humans possess episodic memory (tulving_episodic_1972; tulving2002episodic) enabling us not only to plan and reason to seamlessly operate in the real world, but also to create new or update existing world models by reasoning across multiple experiences (schacter2007remembering; hassabis2007deconstructing).
@@
To illustrate how GSW can help LLMs reason accurately, we evaluate it on the Episodic Memory Benchmark (EpBench) (huet_episodic_2025), that has recently been introduced as a way to benchmark the episodic memory-like capabilities of LLMs. Following are excerpts from two different documents that relate to an entity, Carter Stewart, in this EpBench dataset:
@@
An agent reading the narrative in the first document faces a fundamentally different challenge than traditional fact retrieval. It must understand that “he” refers to a nameless protagonist, who attended a scientific conference where Carter Stewart spoke. The narrator’s spatial context (Metropolitan Museum of Art) and temporal context (September 22, 2026), are stated only indirectly and more importantly have to be also assigned to Carter Stewart who is a presenter. GSW is able to create such representations as part of its working memory construction task: “Carter Stewart: Role: A presenter at a Scientific Conference; Date: September 22, 2026, morning session; Location: The Metropolitan Museum of Art, Topic: statistical analysis; Implements Used: presentation boards and holographic projectors.” The second document is more straightforward and GSW creates a memory trace such as: “Carter Stewart; Role: a researcher and presenter; Location: Bethpage Black Course; Date: March 23, 2024, Did What?: Presented his research findings at a Scientific Conference.” A visualization of the steps of how GSW constructs its working memory is shown in Fig. 6-8 in Appendix B.
@@
When presented with a task such as “List all the unique locations and dates where Carter Stewart made presentations at Scientific Conference events.” a query resolution module (see Section Question Answering with GSW) searches through the GSW constructed from all 200 documents and identifies entities mentioned in the query (e.g., Carter Stewart) that match query’s intent (e.g., a presenter at scientific conference; another entity named Carter Stewart whose role is that of a baker by profession would be ignored) and then returns just the relevant portion of its memory, as in the preceding paragraph. This results in highly targeted and short texts that an LLM has to reason through to provide an answer. In contrast, current structured RAG methods are designed to facilitate retrieval of either whole chunks or community-level summaries that have different levels of similarity to the entities and other phrases in the query. For example, for this query (see Appendix D) GraphRAG’s (Edge2025GraphRAG) summarization missed that Carter Stewart was at the same location as the protagonist in Document #1, and included irrelevant text chunks which led to a list that misses one location and hallucinates two erroneous locations. HippoRAG2 (JimenezGutierrez2025HippoRAG) retrieves the full text of both the relevant documents, along with many other documents, overwhelming the LLM and leading it to hallucinate three erroneous locations. For a more detailed comparison, see Section Concluding Remarks and Limitations, and Tables 1, 4, and 3.
@@
In the rest of this paper, we detail the GSW framework (Section The Generative Semantic Workspace (GSW) Framework) and present a rigorous evaluation on two versions of the EpBench benchmark (Section Question Answering with GSW). The results demonstrate a significant improvement over existing methods. On the EpBench-200 corpus, GSW achieves a state-of-the-art F1-score of 0.85, outperforming strong structured RAG baselines. This advantage is particularly pronounced in the most demanding queries requiring synthesis across as many as 17 different documents, where GSW improves recall by up to 20% over the next best approach as detailed in Table 2. Furthermore, GSW is efficient, reducing the number of context tokens sent to the LLM by 51% compared to the most token-efficient baseline, drastically lowers inference costs and reducing the rate of hallucination in question answering (see Table 3). We further show that this powerful combination of accuracy and token efficiency holds at scale; on the EpBench-2000 corpus, a 10x larger dataset, GSW again achieves a state-of-the-art F1-score of 0.773, outperforming the best baseline by more than 15% on overall recall (Table 4), positioning GSW as a robust and scalable solution for equipping LLMs with effective episodic memory.
@@
Results and discussions are summarized in Section Results and Discussion and a review of related literature is presented Section Related Work. Finally, limitations and future work are discussed in Section Concluding Remarks and Limitations, and a detailed Appendix provides supporting evidence, including manual evaluations performed to validate the power of GSW’s episodic memory capabilities.
@@
The Generative Semantic Workspace (GSW) Framework
In neuroscience, the neocortex is believed to encode hierarchical abstractions of entities, roles, and event templates (george2009towards; Botvinik2008; felleman1991distributed). The hippocampus, especially the CA3 module, plays a complementary role by binding these representations into coherent spatiotemporal sequences (teyler1986hippocampal; rolls2013quantitative; eichenbaum2004hippocampus). During sleep, this neocortical-hippocampal system engages in experience replay, a process through which episodic traces are reactivated in reverse or forward order to consolidate memory and refine internal models (olafsdottir2018role; louie2001temporally; wilson1994reactivation). This back and forth supports both persistence and prediction of memory (mcclelland1995there; rasch2013sleep), key features of episodic memory.
Motivated by this biological architecture (see Fig 1), an effective memory framework requires a structured representation capable of encoding actors along with their evolving roles and states. Crucially, this representation must be capable of spatiotemporal grounding, linking entities and their interactions to specific times and locations, much like the binding function of the hippocampus. Finally, the framework must possess a process for consolidating and updating these structures as new information arrives, mirroring the way the neocortical-hippocampal loop constantly refines its world model. This process of building an evolving model is illustrated with a detailed end-to-end example in Appendix B.
From Episodic Memory to Generative Modeling of Situations and Narratives: The central challenge, therefore, is to create a continuously evolving semantic model, which requires a bidirectional mapping between text and a structured representation. While early symbolic frameworks like PropBank (propbank) and FrameNet (framenet) attempted this, they were not designed for this full bidirectional process, relying instead on fixed ontologies that lacked the necessary probabilistic and dynamic interpretation.
LLMs now make this bidirectional mapping tractable. They can both infer concise semantic identifiers from text and generate coherent narratives from those identifiers. This enables a new, efficient memory model where compact semantic traces are stored and reactivated in context. The formal model is presented next, and its approach is validated in Appendix I, where a human evaluation shows a strong preference for the GSW semantic maps over those from frameworks like PropBank and FrameNet.
A Probabilistic Model for Semantic Memory: The Operator Framework
@@
Enabling Recursive Updates: A State Space Approach (The Reconciler Framework)
@@
Question Answering with GSW
@@
EpBench: An Episodic Memory Benchmark
Our experiments utilize the Episodic Memory Benchmark (EpBench) (huet_episodic_2025), a benchmark specifically designed to evaluate the capabilities of LLMs for episodic memory recall and reasoning over long narratives. Unlike many standard Question Answering (QA) benchmarks (kovcisky2018narrativeqa; zhang_inftybench_2024; yang2018hotpotqa) – focusing on localized factual retrieval – EpBench targets core episodic capabilities: remembering specific events situated in unique spatiotemporal contexts and distinguishing between recurring events involving the same actors.
@@
Table 1: EpBench-200 Dataset Statistics. The statistics for around 1M token Epbench-2000 are presented in Appendix F
EpBench documents are structured as synthetic books generated chapter-by-chapter from event templates (detailing date, location, entity, content) sampled from a larger universe, ensuring recurring elements that necessitate disambiguation and temporal tracking. Chapters are generated via LLM prompts and verified for coherence. Moreover, the same time/location/actors (collectively referred to as cues) appear across multiple chapters. For our evaluation, we use both the standard 200-chapter version and the extended 2000 chapter version of the dataset and report its Statistics in Table 1 Appendix F.
Evaluation Metrics
To evaluate model performance on the EpBench dataset’s queries (detailed in Section EpBench: An Episodic Memory Benchmark), we adopt the LLM-as-a-Judge evaluation paradigm (zheng_judging_2023). For consistency, we strictly follow the LLM-based answer processing and extraction procedure outlined by the EpBench benchmark authors. This approach accounts for the possibility that model responses might be longer or more elaborate than the typically concise ground truth answers. These LLM extracted answers are then used to compute Precision, Recall and F1 scores which we report in Table 2
Baseline Methods
We compare GSW against several baseline approaches: Vanilla LLM, standard Embedding-based RAG (karpukhin-etal-2020-dense; ram_-context_2023) for which we utilized the Voyage-03111https://blog.voyageai.com/2024/09/18/voyage-3/ embedding model selected for its strong performance on retrieval benchmarks (thakur2021beir; muennighoff2022mteb) , and the structured RAG methods GraphRAG (Edge2025GraphRAG), HippoRAG2(JimenezGutierrez2025HippoRAG), and LightRAG (Guo2024LightRAG). We detail the hyper-parameter settings for all baselines in Appendix E .
Implementation Details
The GSW Operator (Section A Probabilistic Model for Semantic Memory: The Operator Framework) and Reconciler (Section Enabling Recursive Updates: A State Space Approach (The Reconciler Framework)) were implemented by prompting GPT-4o (hurst2024gpt) according to task-specific instructions, using temperature set to 0 for deterministic behavior. To ensure fair comparison, we standardized both the maximum context utilization (limited to 17 chapters per query, matching the maximum relevant chapters per query) and the answer generation model (GPT-4o) across all evaluated methods. The complete prompts are provided in Appendix A, and API interactions were managed using the Bespoke Curator library (bespoke_curator), indexing costs are reported in Appendix H. To generate an answer for a given query, we first identify named entities within the query text. These entities are then matched to corresponding nodes within the current GSW memory ($\mathcal{M}_{n}$) using simple string matching. Summaries for the matched entities – aggregated from the GSW structure – are then retrieved and re-ranked based on semantic similarity to the query. The final re-ranked summaries are provided to the LLM to answer the query as illustrated in Figure 2. An end-to-end QA example is provided in Appendix C.

### v2

@@ Abstract
Large Language Models (LLMs) face fundamental challenges in long-context reasoning: many documents exceed their finite context windows, while performance on texts that do fit degrades with sequence length, necessitating their augmentation with external memory frameworks. Current solutions, which have evolved from retrieval using semantic embeddings to more sophisticated structured knowledge graphs representations for improved sense-making and associativity, are tailored for fact-based retrieval and fail to build the space-time-anchored narrative representations required for tracking entities through episodic events. To bridge this gap, we propose the Generative Semantic Workspace (GSW), a neuro-inspired generative memory framework that builds structured, interpretable representations of evolving situations, enabling LLMs to reason over evolving roles, actions, and spatiotemporal contexts. Our framework comprises an Operator, which maps incoming observations to intermediate semantic structures, and a Reconciler, which integrates these into a persistent workspace that enforces temporal, spatial, and logical coherence. On the Episodic Memory Benchmark (EpBench) (Huet et al. 2025) comprising corpora ranging from 100k to 1M tokens in length, GSW outperforms existing RAG based baselines by up to 20%. Furthermore, GSW is highly efficient, reducing query-time context tokens by 51% compared to the next most token-efficient baseline, reducing inference time costs considerably. More broadly, GSW offers a concrete blueprint for endowing LLMs with human-like episodic memory, paving the way for more capable agents that can reason over long horizons. Code is available at https://github.com/roychowdhuryresearch/gsw-memory.
1 Introduction
@@
Large Language Models (LLMs) have transformed natural language understanding, but their ability to reason over long contexts is still limited by finite input windows. Even with token limits in the millions, large document collections can easily exceed these bounds. Performance can also degrade with context length due to phenomena like “context rot” and “lost-in-the-middle” effects (Liu et al. 2023; Hong et al. 2025). A common workaround is Retrieval-Augmented Generation (RAG), which supplements the LLM’s input with only the most relevant retrieved content at query time. Standard RAG pipelines split documents into smaller chunks, encode them into dense embeddings, and retrieve the top-matching chunks based on semantic similarity to the query—allowing the LLM to focus on a relevant subset of the corpus during inference.
@@
A key limitation of standard RAG methods is that each text chunk is embedded independently, which can lead to incomplete retrieval when a query depends on information spread across multiple chunks. Because similarity scores are computed in isolation, essential context may be missed. To address this, more recent approaches have adopted structured representations — such as knowledge graphs — that explicitly model relationships between chunks across the corpus. At query time, these graphs are traversed or queried to retrieve semantically connected chunks, enabling LLMs to perform more effective multi-hop reasoning and question answering (Gutiérrez et al. 2025a, b; Edge et al. 2025; Guo et al. 2025).
@@
These methods have primarily been evaluated on fact-rich documents such as Wikipedia pages (Yang et al. 2018; Ho et al. 2020; Trivedi et al. 2022). Yet the vast majority of texts that LLMs encounter are not lists of facts but narratives of evolving real-world situations. Crime reports, political briefings, corporate filings, legislative records, war dispatches, and multi-day news coverage all describe actors (people, organizations, nations) that adopt roles (suspect, regulator, bidder, combatant) and transition through states (arrested → arraigned → released; startup → unicorn → acquired) while interacting across space and time (Shahsavari et al. 2020a, b; Tangherlini et al. 2020).
@@
We contend that reasoning over such documents would be much more accurate and energy efficient, if one indexed the documents in terms of an internal world model— a structured representation that keeps track of who is involved, what was done, where and when events occur, how roles change, and what consequences follow. Indeed, to achieve such a goal, humans possess episodic memory (Tulving 1972, 2002) enabling us not only to plan and reason to seamlessly operate in the real world, but also to create new or update existing world models by reasoning across multiple experiences (Schacter et al. 2007; Hassabis and Maguire 2007).
@@
To illustrate how GSW can help LLMs reason accurately, we evaluate it on the Episodic Memory Benchmark (EpBench) (Huet et al. 2025), that has recently been introduced as a way to benchmark the episodic memory-like capabilities of LLMs. Following are excerpts from two different documents that relate to an entity, Carter Stewart, in this EpBench dataset:
@@
An agent reading the narrative in the first document faces a fundamentally different challenge than traditional fact retrieval. It must understand that “he” refers to a nameless protagonist, who attended a scientific conference where Carter Stewart spoke. The narrator’s spatial context (Metropolitan Museum of Art) and temporal context (September 22, 2026), are stated only indirectly and more importantly have to be also assigned to Carter Stewart who is a presenter. GSW is able to create such representations as part of its working memory construction task: “Carter Stewart: Role: A presenter at a Scientific Conference; Date: September 22, 2026, morning session; Location: The Metropolitan Museum of Art, Topic: statistical analysis; Implements Used: presentation boards and holographic projectors.” The second document is more straightforward and GSW creates a memory trace such as: “Carter Stewart; Role: a researcher and presenter; Location: Bethpage Black Course; Date: March 23, 2024, Did What?: Presented his research findings at a Scientific Conference.”
@@
When presented with a task such as “List all the unique locations and dates where Carter Stewart made presentations at Scientific Conference events.” a query resolution module (see Section 3) searches through the GSW constructed from all 200 documents and identifies entities mentioned in the query (e.g., Carter Stewart) that match query’s intent (e.g., a presenter at scientific conference; another entity named Carter Stewart whose role is that of a baker by profession would be ignored) and then returns just the relevant portion of its memory, as in the preceding paragraph. This results in highly targeted and short texts that an LLM has to reason through to provide an answer. In contrast, current structured RAG methods are designed to facilitate retrieval of either whole chunks or community-level summaries that have different levels of similarity to the entities and other phrases in the query. For example, for this query GraphRAG’s (Edge et al. 2025) summarization missed that Carter Stewart was at the same location as the protagonist in Document #1, and included irrelevant text chunks which led to a list that misses one location and hallucinates two erroneous locations. HippoRAG2 (Gutiérrez et al. 2025b) retrieves the full text of both the relevant documents, along with many other documents, overwhelming the LLM and leading it to hallucinate three erroneous locations. For a more detailed comparison, see Section 6, and Tables 1, 4, and 3.
@@
In the rest of this paper, we detail the GSW framework (Section 2) and present a rigorous evaluation on two versions of the EpBench benchmark (Section 3). The results demonstrate a significant improvement over existing methods. On the EpBench-200 corpus, GSW achieves a state-of-the-art F1-score of 0.85, outperforming strong structured RAG baselines. This advantage is particularly pronounced in the most demanding queries requiring synthesis across as many as 17 different documents, where GSW improves recall by up to 20% over the next best approach as detailed in Table 2. Furthermore, GSW is efficient, reducing the number of context tokens sent to the LLM by 51% compared to the most token-efficient baseline, drastically lowers inference costs and reducing the rate of hallucination in question answering (see Table 3). We further show that this powerful combination of accuracy and token efficiency holds at scale; on the EpBench-2000 corpus, a 10x larger dataset, GSW again achieves a state-of-the-art F1-score of 0.773, outperforming the best baseline by more than 15% on overall recall (Table 4), positioning GSW as a robust and scalable solution for equipping LLMs with effective episodic memory.
@@
Results and discussions are summarized in Section 4 and a review of related literature is presented in Section 5. Finally, limitations and future work are discussed in Section 6. The full version of the paper  [^1] The full version of this paper with a detailed technical appendix and clear visualization of the GSW using an example is available at https://arxiv.org/abs/2511.07587. provides supporting evidence, including manual evaluations performed to validate the power of GSW’s episodic memory capabilities.
@@
2 The Generative Semantic Workspace (GSW) Framework
In neuroscience, the neocortex is believed to encode hierarchical abstractions of entities, roles, and event templates (George and Hawkins 2009; Botvinick 2008; Felleman and Van Essen 1991). The hippocampus, especially the CA3 module, plays a complementary role by binding these representations into coherent spatiotemporal sequences (Teyler and DiScenna 1986; Rolls 2013; Eichenbaum 2004). During sleep, this neocortical-hippocampal system engages in experience replay, a process through which episodic traces are reactivated in reverse or forward order to consolidate memory and refine internal models (Ólafsdóttir et al. 2018; Louie and Wilson 2001; Wilson and McNaughton 1994). This back and forth supports both persistence and prediction of memory (McClelland et al. 1995; Rasch and Born 2013), key features of episodic memory.
Motivated by this biological architecture (see Fig 1), an effective memory framework requires a structured representation capable of encoding actors along with their evolving roles and states. Crucially, this representation must be capable of spatiotemporal grounding, linking entities and their interactions to specific times and locations, much like the binding function of the hippocampus. Finally, the framework must possess a process for consolidating and updating these structures as new information arrives, mirroring the way the neocortical-hippocampal loop constantly refines its world model.
From Episodic Memory to Generative Modeling of Situations and Narratives: The central challenge, therefore, is to create a continuously evolving semantic model, which requires a bidirectional mapping between text and a structured representation. While early symbolic frameworks like PropBank (Kingsbury and Palmer 2002) and FrameNet (Baker et al. 1998) attempted this, they were not designed for this full bidirectional process, relying instead on fixed ontologies that lacked the necessary probabilistic and dynamic interpretation.
LLMs now make this bidirectional mapping tractable. They can both infer concise semantic identifiers from text and generate coherent narratives from those identifiers. This enables a new, efficient memory model where compact semantic traces are stored and reactivated in context. The formal model is presented next.
2.1 A Probabilistic Model for Semantic Memory: The Operator Framework
@@
2.2 Enabling Recursive Updates: A State Space Approach (The Reconciler Framework)
@@
3 Question Answering with GSW
@@
3.1 EpBench: An Episodic Memory Benchmark
Our experiments utilize the Episodic Memory Benchmark (EpBench) (Huet et al. 2025), a benchmark specifically designed to evaluate the capabilities of LLMs for episodic memory recall and reasoning over long narratives. Unlike many standard Question Answering (QA) benchmarks (Kočiskỳ et al. 2018; Zhang et al. 2024; Yang et al. 2018) – focusing on localized factual retrieval – EpBench targets core episodic capabilities: remembering specific events situated in unique spatiotemporal contexts and distinguishing between recurring events involving the same actors (Holur et al. 2023, 2022).
@@ 3.1 EpBench: An Episodic Memory Benchmark
Table 1: EpBench-200 Dataset Statistics.
EpBench documents are structured as synthetic books generated chapter-by-chapter from event templates (detailing date, location, entity, content) sampled from a larger universe, ensuring recurring elements that necessitate disambiguation and temporal tracking. Chapters are generated via LLM prompts and verified for coherence. Moreover, the same time/location/actors (collectively referred to as cues) appear across multiple chapters. For our evaluation, we use both the standard 200-chapter version and the extended 2000 chapter version of the dataset and report its Statistics in Table 1.
3.2 Evaluation Metrics
To evaluate model performance on the EpBench dataset’s queries (detailed in Section 3.1), we adopt the LLM-as-a-Judge evaluation paradigm (Zheng et al. 2023). For consistency, we strictly follow the LLM-based answer processing and extraction procedure outlined by the EpBench benchmark authors. This approach accounts for the possibility that model responses might be longer or more elaborate than the typically concise ground truth answers. These LLM extracted answers are then used to compute Precision, Recall and F1 scores which we report in Table 2
3.3 Baseline Methods
We compare GSW against several baseline approaches: Vanilla LLM, standard Embedding-based RAG (Karpukhin et al. 2020; Ram et al. 2023) for which we utilized the Voyage-03222https://blog.voyageai.com/2024/09/18/voyage-3/ embedding model selected for its strong performance on retrieval benchmarks (Thakur et al. 2021) , and the structured RAG methods GraphRAG (Edge et al. 2025), HippoRAG2(Gutiérrez et al. 2025b), and LightRAG (Guo et al. 2025).
3.4 Implementation Details
The GSW Operator (Section 2.1) and Reconciler (Section 2.2) were implemented by prompting GPT-4o (Hurst et al. 2024) according to task-specific instructions, using temperature set to 0 for deterministic behavior. To ensure fair comparison, we standardized both the maximum context utilization (limited to 17 chapters per query, matching the maximum relevant chapters per query) and the answer generation model (GPT-4o) across all evaluated methods. To generate an answer for a given query, we first identify named entities within the query text. These entities are then matched to corresponding nodes within the current GSW memory ($\mathcal{M}_{n}$) using simple string matching. Summaries for the matched entities – aggregated from the GSW structure – are then retrieved and re-ranked based on semantic similarity to the query. The final re-ranked summaries are provided to the LLM to answer the query as illustrated in Figure 2.